OpenStudy (anonymous):

Keisha's parents want to save $20,000 in her college savings account over the next 15 years. They have $8,000 to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth

OpenStudy (anonymous):

20,000 = 8,000 + 8000 * r * 15 r = (20000 - 8000) / (8000* 15) = 0.10 10% its simple interest not compound

OpenStudy (anonymous):

Denebel i think your wrong because it has nothing in the multiple choices related to 6.29